Water is flowing in a fire hose with a velocity of 1 m/s and a pressure of 200000 Pa. At the nozzle the pressure decreases to atmospheric pressure (101300 Pa), there is no change in height. Using Bernoulli equation, the velocity of the water exiting the nozzle is(Hint: The density of water is 1000 kg/m3 and gravity g is 9.8 m/s2. Pay attention to units!)]


1
Expert's answer
2021-03-29T08:58:08-0400
"P_1+0.5\rho v_1^2=P_2+0.5\rho v_2^2\\200000+0.5(1000)(1)^2=101300+0.5(1000)v_2^2\\v_2=14\frac{m}{s}"